Longitudinal waves are set up in an aluminum rod. Calculate the velocity of the wave if the modulus of elasticity of aluminum is 6.9×10^11 dynes per sq. cm. and a density of the aluminum is 2.69 gram per cu. cm.


1
Expert's answer
2021-05-10T02:42:45-0400

1 dyn/cm² to N/m² = 0.1 N/m²

6.9×10"^{11}" dynes per sq. cm= "6.9 \times 10^{10} N/m^2"


density of aluminum = 2.69 g/cm"^3" = 2690 "Kg\/m^3"


Velocity of a wave in solid medium "\nu= \sqrt{\frac{Young's modulus}{density of material}}"

"\nu= \sqrt{\frac{Y}{\rho}}= \sqrt{\frac{6.9 \times 10^{10}}{2690}} = 5064 m/s"